OpenStudy (anonymous):

farmer ed has 2,000 meters of fencing, and wants to enclose a rectangular plot that borders on a river. If farmer ed does not fence the side along the river, what is the largest area that can be enclosed?

OpenStudy (campbell_st):

well start by drawing a picture  so the perimeter is P = 2x + l since there is 2000 meters of fencing its 2000 = 2x + l or making l the subject l = 2000 - 2x now the Area = length * width or after substituting the length value from the perimeter A = (2000 -2x) * x which is \[A = 2000x - 2x^2\] so all you need to do now is 1. find the 1st derivative 2. solve the derivative for x and you'll have the width that gives the max area, substitute it into the perimeter equation for the the length that gives the max area. hope this helps